OpenStudy (anonymous):

To discourage random guessing on a multiple choice exam the professor assigns 5 points for a correct response deducts 2 points for an incorrect answer and gives 0 points for leaving the question blank what is your score on a 35 question exam if you have 22 correct and 7 incorrect answers?

OpenStudy (anonymous):

Express your score as a percentage of the total points possible

OpenStudy (misty1212):

HI!! \[32\times 5-7\times 2\] is the raw score then divide by 35 to get a percent i guess

OpenStudy (anonymous):

Where did you get 32 from?

OpenStudy (misty1212):

it was a mistake \[22\times 5-7\times 2\]